OpenStudy (anonymous):

{(x,y): x-3y=4} x-3y=4 2x-6y=Q Find Q? I'm sooo confused.

OpenStudy (anonymous):

You are given x,y such that (the colon there means such that) x-3y=4. In other words for any x and y youo choose they must satisfy that condition. so if you set x = 2 then you need y such that 2-3y=4 or rearranged (2-4)/(-3) = y. Does this help?

OpenStudy (anonymous):

Q = 8?

OpenStudy (anonymous):

Yep! It is just the first equation multiplied by 2. Good job!